[Ncep.nhc.nco_contacts] TSB Morning Rounds - Wed Jul 05, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Wed Jul 5 11:52:52 UTC 2017


Here is today's summary of NHC computer operations:
--- Mark

----------

TAFB:
The TAFB shift log has not been working properly since Friday, so today's  
morning rounds includes items from the last few days. Monica will look into  
the shift log problem today.

There was a brief interruption of some GOES-W sectors in N-AWIPS on  
Tuesday. This appears to have been caused by problems upstream.

Lightning data has stopped coming in to N-AWIPS since Sunday. TSB will look  
into this.

The Shiner and Grunt N-AWIPS systems are having trouble printing. Also,  
shiner has been freezing up. Mark will put in an NCO trouble ticket to  
address these problems.

There was a GFE issue on Monday. TAFB typed in the preliminary track for  
AL94 to create a "PRETCM" to set up the grids, but the TCM WIND TOOL would  
not read the PRETCM. Chris Mello will follow up.


HSU:
No new issues.
Other items:

NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170705/0631a1d4/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list